OpenStudy (aripotta):

If j, k, and n are consecutive integers such that 0 < j < k < n and the units (ones) digit of the product jn is 9, what is the units digit of k? a. 0 b. 1 c. 2 d. 3 e. 4

OpenStudy (nincompoop):

it would be good to start with 3

OpenStudy (aripotta):

i'm confused with units digit?? what do they mean by that

OpenStudy (freckles):

hmmm... so we have j=i-1 k=i n=i+1 where i is integer jn=i^2-1 and jn has number representation: \[c_n10^{n}+c_{n-1}10^{n-1}+c_{n-2}10^{n-2}+\cdots+c_310^3+c_210^2+c_110^1+c_010^0\] and we given c_0 is 9... ... still thinking on how to proceed with this

OpenStudy (nincompoop):

most likely, the units here are referred to the place value 123, to hundreds 12, tens 1, ones

OpenStudy (nincompoop):

0.1 tenth 0.34 hundreth etc.

OpenStudy (nincompoop):

http://www.mathsisfun.com/place-value.html

OpenStudy (aripotta):

i know the answer is a, i just don't know how they got it

OpenStudy (nincompoop):

it cannot be a since 0 < j 0 = 0 and not 0 < 0

OpenStudy (freckles):

\[i^2-1=.....9 \\ \] what is the unit digit for i^2 if you all 1 on both sides

OpenStudy (aripotta):

that's what the answer key says though, nin

OpenStudy (nincompoop):

oh LAUGHING OUT LOUD well then

OpenStudy (freckles):

j>0 doesn't mean the units digit of j is 0

OpenStudy (freckles):

j could be 10 for all we know 10 is still bigger than 0

OpenStudy (freckles):

I think the easier way it to add 9 on both sides of i^2-1=.....9

OpenStudy (aripotta):

i'm like, really confused right now

OpenStudy (nincompoop):

so how do we satisfy the product jn = 9

OpenStudy (freckles):

and say that i^2 has unit digit?

OpenStudy (nincompoop):

and j < n

OpenStudy (freckles):

i^2-1=.....9 i^2=.......0 so i has to to have unit digit?

OpenStudy (freckles):

i will let you finish it from here

OpenStudy (nincompoop):

I do not see why j = 0 maybe because I just woke up. I will think about this, but right now sorry that I am not of any help

OpenStudy (aripotta):

freckles, i don't know what to do ._.

OpenStudy (freckles):

first do you understand what i have written

OpenStudy (aripotta):

no :x

OpenStudy (nincompoop):

he is suggesting that the values j and k are not straight-forward so he is indicating them as i j = i - 1 k = i n = i + 1 which has to satisfies j < k <n but our condition is 0 < j < k < n this means that our j cannot be 0

OpenStudy (freckles):

jn is i^-1 and jn has one vale 9

OpenStudy (freckles):

i^2-1*

OpenStudy (nincompoop):

is that i^{2-1} ?

OpenStudy (freckles):

so jn looks like some numbers then 9

OpenStudy (freckles):

no (i-1)(i+1)=i^2-1

OpenStudy (freckles):

i^2-1=......9

OpenStudy (nincompoop):

are you referring to THE imaginary number @freckles ?

OpenStudy (aripotta):

ok i get that part now

OpenStudy (freckles):

those dots just represent the digits in the 10s 100s and so on place value

OpenStudy (freckles):

lol no i said i was an integer above

OpenStudy (nincompoop):

then use Z next time LAUGHING OUT LOUD

OpenStudy (freckles):

are you purposely giving me a hard time :p

OpenStudy (freckles):

look i defined i above

OpenStudy (freckles):

in my first post

OpenStudy (aripotta):

freckles, i get you lol. what we do from here?

OpenStudy (aripotta):

what do we do*

OpenStudy (freckles):

we have i^2-1=.......9

OpenStudy (freckles):

if you add 1 to both sides that changes the unit digit only for the right hand side

OpenStudy (freckles):

after all 1+9=10 so you i^2=.......0

OpenStudy (freckles):

so i=.......0

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{0<j<k<n\\~\\ \implies 0<j<j+1<j+2\\~\\ (j\times n )mod 10\equiv 9\\~\\ (j\times (j+2) )mod 10\equiv 9\\~\\ j=19\\~\\ 0<j<k<n\\~\\ \implies 0<19<20<21\\~\\ j\times n=19\times 21=399\\ \text{it satisfies the condition}\\~\\ \text{units digit of k is zero}\\ }\end{align}\)

OpenStudy (aripotta):

wait whaaaat

OpenStudy (aripotta):

i still don't get it. i guess i'm stupid

OpenStudy (mathmath333):

u can use the modular arthmatic method if u know it

OpenStudy (aripotta):

i do not know it :/

OpenStudy (freckles):

the only integer number i can square and get 0 as a unit digit is a number that also has unit value 0 right? 0^2=0 1^2=1 2^2=4 3^2=9 4^2=16 5^2=25 6^2=36 ... 10^2=100

OpenStudy (aripotta):

ok

OpenStudy (freckles):

so if i^2=.......0 then i=..........0

OpenStudy (aripotta):

but how do you get the answer of 0?

OpenStudy (freckles):

we just did

OpenStudy (freckles):

recall i is k

OpenStudy (aripotta):

oh wait i think i get it now. so because i^2 - 1 = ....9 and adding 1 to both sides would be i^2 = .....0 and you can only get 0 in the unit place when you square a number, only if the number itself ends in 0, then the answer must be 0

OpenStudy (freckles):

yeah

OpenStudy (aripotta):

ok thank you very much :)